# Green Functions in Small Characteristic

### v2

@@ 4 Green functions by the Lusztig-Shoji algorithm
Let $\chi\in\textrm{Irr}(W)$ be an $F$-stable irreducible character of $W$. Writing $F$ also for the automorphism of $W$ induced by the Frobenius endomorphism, we consider the semidirect product $W\rtimes\langle F\rangle$; the $W$-conjugacy classes in the coset $WF$ are in bijection with the $F$-conjugacy classes of $W$. The character $\chi$ can be extended to an irreducible character $\tilde{\chi}$ of $W\rtimes\langle F\rangle$. For any character $\tilde{\chi}$ of $W\rtimes\langle F\rangle$ we will consider the linear combination
$Q_{\tilde{\chi}}:=\sum_{i=1}^{k}\frac{1}{|C_{W}(w_{i}F)|}\tilde{\chi}(w_{i}F)Q_{w_{i}}.$
@@ 5.1 Permutation characters by Deligne-Lusztig characters
So, its restriction to unipotent elements is $Q_{\bar{\chi}}$ for the trivial character $\bar{\chi}$ on $WF$.
@@
This shows that the restriction to unipotent elements is $Q_{\bar{\chi}}$ where $\bar{\chi}$ is the permutation character of $WF$ on $W_{J}F\subseteq WF$.
@@ 6 Application to $\mathbf{E_{8}(q)}$
Now we use Section 5.3 and compute the values of this permutation character on the elements $z_{52},z_{53},z_{54}$ given in Mizuno [Miz80] and find the first of the tuples above which corresponds to $\zeta_{u,-1}=-1$ and $\zeta_{u,\rho}=1$. This is the exception stated in our Theorem 6.1. Note that by Proposition 4.1 this exception only occurs when $q$ is an odd power of $2$ (or $q\equiv-1(3)$).
Remarks: The computation of each value took about 5 seconds, there are $3628800$ $J$-reduced elements $w$, but only about $500$ of them do not fulfill the criterion in Section 5.5. Note that with our computation we can actually show that the elements $z_{52},z_{53},z_{54}$ in Mizuno’s list are indeed in the class $C=D_{8}(a_{3})$ because the same values do not occur on other classes. In fact, we found an error in Mizuno’s list by computing the value of $z_{44}$. This also yields the value $177889$ and this shows that this element is conjugate to $z_{53}$.
@@
As in step (6) we can use the parabolic subgroup for $J$ of type $D_{4}$. The four possibilities for the two unknown scalars always lead to the same four character values, but they differ in which of the four values appears twice. For $q=2$ the four values are $\{6785,2625,6401,2241\}$. Using the five representatives $z_{i}$, $27\leq i\leq 31$, in Mizuno’s list and computing the character values for these elements we get the value $6785$ twice, namely for $z_{27}$ and $z_{31}$. This shows that also the scalars $\zeta_{u,\epsilon^{\prime}}=\zeta_{u,\epsilon^{\prime\prime}}=1$.

### v3

@@ 4 Green functions by the Lusztig-Shoji algorithm
Let $\chi\in\textrm{Irr}(W)$ be an $F$-stable irreducible character of $W$. Writing $F$ also for the automorphism of $W$ induced by the Frobenius endomorphism, we consider the semidirect product $W\rtimes\langle F\rangle$; the $W$-conjugacy classes in the coset $WF\subseteq W\rtimes\langle F\rangle$ (the $W$-coset containing $F$) are in bijection with the $F$-conjugacy classes of $W$. The character $\chi$ can be extended to an irreducible character $\tilde{\chi}$ of $W\rtimes\langle F\rangle$.
For any class function $\tilde{\chi}$ on the $W$-conjugacy classes of the coset $WF$ we define the linear combination
$Q_{\tilde{\chi}}:=\frac{1}{|W|}\sum_{w\in W}\tilde{\chi}(wF)Q_{w}=\sum_{i=1}^{k}\frac{1}{|C_{W}(w_{i}F)|}\tilde{\chi}(w_{i}F)Q_{w_{i}}.$
@@ 5.1 Permutation characters by Deligne-Lusztig characters
So, its restriction to unipotent elements is $Q_{\bar{\chi}}$ for the trivial character $\bar{\chi}$ of the coset $WF$ (see the definition of $Q_{\tilde{\chi}}$ in Section 4).
@@
This shows that the restriction to unipotent elements is $Q_{\bar{\chi}}$ where $\bar{\chi}$ is the permutation character of $WF$ on $W_{J}F\subseteq WF$ (defined as $\bar{\chi}(wF)=1/|W|\cdot\sum_{x\in W,\;(wF)^{x}\in W_{J}F}1$, which specializes to ${1_{W_{J}}}^{W}$ in case $F=\textrm{id}_{W}$).
@@ 6 Application to $\mathbf{E_{8}(q)}$
Now we use Section 5.3 and compute the values of this permutation character on the elements $z_{77},z_{78},z_{79}$ given in Mizuno [Miz80] and find the first of the tuples above which corresponds to $\zeta_{u,-1}=-1$ and $\zeta_{u,\rho}=1$. This is the exception stated in our Theorem 6.1. Note that by Proposition 4.1 this exception only occurs when $q$ is an odd power of $2$ (or $q\equiv-1(3)$).
Remarks: The computation of each value took about 5 seconds, there are $3628800$ $J$-reduced elements $w$, but only about $500$ of them do not fulfill the criterion in Section 5.5. Note that with our computation we can actually show that the elements $z_{77},z_{78},z_{79}$ in Mizuno’s list are indeed in the class $C=D_{8}(a_{3})$ because the same values do not occur on other classes (and we see that $F$ acts trivially on the component group of $z_{77}$). In fact, we found an error in Mizuno’s list by computing the value of $z_{64}$. This also yields the value $177889$ and this shows that this element is conjugate to $z_{78}$.
@@
As in step (6) we can use the parabolic subgroup for $J$ of type $D_{4}$. The four possibilities for the two unknown scalars always lead to the same four character values, but they differ in which of the four values appears twice. For $q=2$ the four values are $\{6785,2625,6401,2241\}$. Using the five representatives $z_{i}$, $i\in\{44,46,47,48,49\}$, in Mizuno’s list and computing the character values for these elements we get the value $6785$ twice, namely for $z_{44}$ and $z_{49}$. This shows that also the scalars $\zeta_{u,\epsilon^{\prime}}=\zeta_{u,\epsilon^{\prime\prime}}=1$.